OpenStudy (anonymous):

2 boxes of pens and 6 boxes of pencils cost $14.00. 4 boxes of pens and 3 boxes of pencils cost 14.50. What is the cost of each box of pencils? Come up with a system of equations to solve.

OpenStudy (katrinakaif):

Using elimination first can help rather than substitution

OpenStudy (katrinakaif):

Lets multiply the first equation by -3 giving us -3(2b + 6p) = (14 * (-3) -6b - 18p = 42 Now the second equation by 2 2 (3b + 4p) = 14.50 (2) 6b + 8p = 29

OpenStudy (anonymous):

Okay. This is a classic example of simultaneous equations. We're using 'i' for (ink, to represent pens), and 'g' for graphite, to represent pencils. 2i + 6g = 14 That's your first sentence in mathematical form. Now, we put the next one into graphical form as well. 4i + 3g = 14.5 So, we'll half the top equation and find out what one box of pens is equal to. i + 3g = 7. Let's find out what pens alone are equal to by rearranging it. i = 7 - 3g. Now, we can put this into the other equation, if we know the price of pens. 4(7 - 3g) + 3g = 14.5 28 - 12g + 3g = 14.5 28 - 9g = 14.5 Rearrange it for 13.5 = 9g Divide each side by 9, and we have one box of pencils costing 1.5. Now we know that g = 1.5, we can put it back into the original equation. 2i + 6g = 14 2i + 6*1.5 = 14 2i + 9 = 14 2i = 5 i = 2.5 SO, a box of pens costs $2.50, and a box of pencils costs $1.50